OpenStudy (anonymous):

(x+7)^(2/3) = 9 x can equal ?

OpenStudy (anonymous):

\[(x+7)^\frac23=9\]

sam (.sam.):

\[\Huge(x+7)^\frac23=9\] \[\Huge(x+7)=9^\frac32\] \[\Huge x=9^\frac32-7\]

OpenStudy (anonymous):

x is supposed to have 2 possible answers and somehow i got 20...

OpenStudy (anonymous):

i dont know how to show it becasue im bad with the equation maker thing :(

OpenStudy (anonymous):

Is the problem expression the following?\[(x+7)^{2/3}=9 x \]

sam (.sam.):

\[\Huge(x+7)^\frac23=9\] \[\Huge(x+7)^2=9^{3}\] \[\Huge(x^2+14x+49)=9^{3}\] \[\Huge(x^2+14x-680)=0\] (x+34)(x-20)=0

sam (.sam.):

2 solutions there

OpenStudy (anonymous):

ok , thanks so -34 and 20?

sam (.sam.):

yes

OpenStudy (anonymous):

thanks :D

sam (.sam.):

welcome ;)